# Probing the Origins I: Generalised Additive Model inference of birth radii for Milky Way stars in the solar vicinitythanks: Catalogue reference

### v1

@@ 3.2.3 Other signs of migration
In the right panel, we greatly enlarged the marker sizes to make the high $\langle e\rangle$ region more prominent. As a result, stars with the largest eccentricities — least likely to be affected by churning — cluster on the left side of the plot (e.g. Khoperskov et al., 2020). These stars show a wider range of metallicities, smaller $\langle R_{\rm{g}}\rangle$ values (closer to the Galactic centre), and tend to be older. Conversely, in the left panel, stars with the largest radial differences, $\Delta R$, span a range of $\langle R_{\rm{g}}\rangle$ values but are predominantly metal-poor, with ages ranging from young to intermediate (about 0 to 4–5 Gyr). Although the marker size indicates stars with larger variation in their past and current orbital radii, which can be a sign of radial migration, it does not necessarily confirm it, since we are not accounting for errors-in-measurements here. This will be further discussed in Sect. 3.3.
@@
3.3 Churned vs. blurred/undisturbed stars
@@ 3.3.2 Angular momentum features of motion-stratified stellar populations
Figure 18: Median angular momenta in the $r$, $\phi$, and $z$ directions – $\langle L_{r}\rangle$, $\langle L_{\phi}\rangle$, $\langle L_{z}\rangle$ – plotted against the median metallicity, $\langle\rm{[Fe/H]}\rangle$. Each panel presents the angular momentum in a specific direction. Marker colours represent distinct HC groups, consistent with the other figures in this paper, while marker shapes indicate different stellar movements: inward churning, outward churning, and blurring (labelled as “equal”). The lines serve as visual guides rather than regressions, highlighting potential trends: the solid line connects HC groups with inward churning, the dashed line corresponds to outward churning, and the dotted line represents stars with unchanged radii (undisturbed or blurred).
In Fig. 18, we display $\langle L_{r}\rangle$, $\langle L_{\phi}\rangle$, and $\langle L_{z}\rangle$ against $\langle\rm{[Fe/H]}\rangle$, [Fe/H] being a rough proxy for $\langle R_{\rm{b}}\rangle$ as illustrated by the radial metallicity gradients shown in this paper. The colours in this figure align with those used throughout this paper, representing the different HC metallicity-stratified groups. Marker shapes indicate the type of movement (inward/outward-churned; and blurred/undisturbed, tagged as “equal”), with distinct lines connecting markers of the same movement to guide the eye towards potential trends (without performing a formal regression analysis).
@@ 3.3.4 $\langle R_{\rm{b}}\rangle$, $\langle R_{\rm{g}}\rangle$, and the OLR
Building on the discussion developed right before the beginning of Sect. 3.3.1, the bar structure likely influences stars formed in the inner regions of the MW, whereas stars moving inward may be more sensitive to the transient spiral arms than to the bar. Although there is ongoing debate regarding the length of the MW’s bar and its components (such as the so-called long-bar; López-Corredoira et al., 2007; Wegg & Gerhard, 2013; Portail et al., 2017; Lucey et al., 2023), recent estimates place its length at $\approx$ 3.5 kpc (Lucey et al., 2023), which could constrain the range of the bar’s influence, particularly in its proximity. These estimates agree with predictions from Khoperskov et al. (2020), which indicate that the bar—and more specifically, its deceleration—contributes to the churning of stars from the inner Galaxy outward toward the OLR.
@@
Figure 19: Same as Fig. 18, but for the action components ($\langle J_{r}\rangle$, $\langle J_{\phi}\rangle$, $\langle J_{z}\rangle$), instead of angular momentum.
Figure 20: $\langle R_{\rm{b}}\rangle$ against $\langle R_{\rm{g}}\rangle$ for all the metallicity- and motion-stratified groups of our sample. The general features of this plot are consistent with those observed in Figs. 18 and 19. The expected outer Lindblad resonance (OLR) radius (estimated by Dehnen, 2000) is marked in yellow.
@@ 3.4 Determining the probable birth radius of the Sun
Here, we investigate what our GAM model estimates as the most probable location of the Sun’s $R_{\rm b}$. As input parameters, we adopted the Sun’s age, $t_{\odot}=4.775\pm 0.039$ Gyr, as provided by Bonanno & Fröhlich (2015). Specifically, we selected the final estimate from Table 3 of Bonanno & Fröhlich (2015), which the authors regard as the most reliable. This estimate was chosen based on the highest Bayes factor, derived using the Irwin+AdelR+Asplund model. For the Sun’s metallicity, we followed the convention [Fe/H]=0, but allowed for an error margin of 0.03 dex, as suggested by the non-local thermal equilibrium (NLTE) estimate presented by Korn et al. (2003) and further discussed in Asplund et al. (2009).
@@ 4 Summary and conclusions
In this study, we harnessed the power of a Generalised Additive Model (GAM) to refine the spatial and chemical grids of established chemical enrichment models for the Milky Way (MW; Magrini et al., 2009). Moreover, this GAM framework was adeptly employed to estimate the birth radii of a select group of Gaia-ESO stars, encompassing a diverse range of metallicities, spanning from metal-poor to super-metal-rich, previously classified into six metallicity-stratified groups through Hierarchical Clustering (HC). Detailed elucidation of our data curation and processing methodologies can be found in the comprehensive exposition by Dantas et al. (2023). Our overarching aim revolved around a quantitative and qualitative exploration of the dynamic effects manifesting within these stars, notably the phenomena of blurring and churning, the latter also known as radial migration. Our conclusions are as follows.
@@
    The GAM has proven to be a useful method for directly deriving the Galactocentric birth radii, $\langle R_{\rm{b}}\rangle$, of the stars in our sample by employing a minimalist approach, using only [Fe/H] and the age of the Universe at the time of the stars’ birth ($t$). This simplicity was intentional, as introducing other available abundances – such as $\alpha$-elements, which are highly correlated with one another and with [Fe/H] – would not necessarily improve the GAM’s performance. Instead, the GAM’s effectiveness is more likely to improve if the underlying chemical enrichment models incorporate a broader range of independent information, such as neutron-capture element abundances, or better reflect the complexities of the MW’s evolutionary history, including multiple gas infall events. Additionally, models that provide a richer dataset with larger age and metallicity bins, as well as smoother radial gradients, would allow for more precise predictions. Conversely, including degenerate information does not enhance the performance of the GAM, further justifying our choice of a minimalist approach in this study.
@@
    We estimated the expected $\langle R_{\rm{b}}\rangle$ for the stars of our sample, using chemical evolution models by Magrini et al. (2009), which were used to run a GAM to expand the reach of the original theoretical models. By comparing their $\langle R_{\rm{b}}\rangle$ with current guiding radii ($\langle R_{\rm{g}}\rangle$), we classified the stars of our sample according to their motion: inward churning, outward churning, and equal motion (undisturbed or blurred). As a consequence, we unveiled a significant migration pattern within the Galaxy:
@@
      We identify a total of about 3/4 of our sample having sustained churning, whereas the remaining $\sim$ 1/4 has either been undisturbed or subject to blurring.
@@
      The most metal-poor stars, in contrast, appear to have originated at larger Galactocentric distances and migrated inward toward the Galaxy’s central regions (approximately 43%), while the largest portion seems to have retained their $\langle R_{\rm{b}}\rangle$ (around 48%). It is important to note, however, that this group may still contain a small number of thick disc interlopers with metallicities that overlap with the thin disc. Despite this possible contamination, the general patterns in their motion remain consistent within this subset.
@@
      Stars with intermediate metallicities exhibit a transitional behaviour, with some migrating outward and others inward. In other words, as median metallicities decrease throughout our metallicity-stratified groups, the proportion of stars migrating outward decreases and those migrating inward increase, as well as those with orbits consistent with their $\langle R_{\rm{b}}\rangle$. This smooth transition is consistent throughout the groups.
@@
    There is a significant age gap between stars influenced by inward and outward churning – for instance, in Group 1 (the second most metal-rich group), this gap is approximately 5 Gyr, and it tends to persist or widen in other metallicity-stratified groups. Stars affected that have kept their orbital radii, on the other hand, exhibit intermediate ages. This age pattern remains consistent across all HC groups.
@@
    Due to the overall old ages of our blurred/undisturbed group, we doubt that most of these stars are undisturbed, since we believe that time is an important factor that dictates the interactions between the stars and the Galactic structures. In other words, we reckon that most stars seem to have undergone blurring.
@@
    Although eccentricity is expected to play a relevant role for churned stars, we do not see any strong relation between eccentricity and type of movement (we do see that stars orbiting at lower $\langle R_{\rm{g}}\rangle$ have slightly larger eccentricities). The eccentricity for all the metallicity- and motion-stratified groups of stars in the sample is somewhat low, with median values ranging from 0.13 to 0.15. There is no observable trend across our HC groups.
@@
    The angular momentum ($L$) analysis results provide valuable insights into the processes that shape stellar motion within our sample. Notably, $\langle L_{z}\rangle$ reveals distinct trends, contrasting with the lack of significant patterns in $\langle L_{r}\rangle$ and $\langle L_{\phi}\rangle$. Across the sample, $\langle L_{z}\rangle$ consistently increases as metallicity decreases, irrespective of stellar motion type. This trend may reflect the chemical enrichment history of the MW, as $L_{z}$ traces patterns within the Galactic disc. Further, age differences between motion-stratified groups strengthen this interpretation. Outward-churned stars, for instance, tend to have higher $L_{z}$ and are systematically older within each metallicity-defined group than their blurred/undisturbed and inward-churned counterparts. Inward-churned stars, conversely, are consistently the youngest, exhibiting lower $L_{z}$ than their blurred/undisturbed peers, who hold intermediate positions in age and $L_{z}$. Another perspective (or interpretation) on the behaviour of $L_{z}$ considers the role of Galactic structures: outward churning and elevated $\langle L_{z}\rangle$ might result from interactions with these structures, whereas inward-churned stars, being younger, appear to “lose” $\langle L_{z}\rangle$ as they move inward. These findings underscore the possible influence of Galactic structures on stellar motion types, particularly concerning $L_{z}$. However, the degree to which these ages align with timescales required for inward or outward churning remains an open question.
@@
    The results observed in our analysis of the action components – $\langle J_{r}\rangle$, $\langle J_{\phi}\rangle$, $\langle J_{z}\rangle$ – corroborate findings from other parameters and reinforce the interpretations discussed. Specifically, $\langle J_{r}\rangle$  which quantifies radial orbital excursions as an integrated quantity, provides key insights into radial orbital stability. Our analysis suggests that stars in Group 6 may have stabilised their orbits earlier. A reasonable hypothesis is their proximity at birth to the outer Lindblad resonance (OLR) radius. This resonance, known to influence orbital dynamics, could have accelerated the damping of radial oscillations, allowing these stars to achieve orbital stability earlier compared to stars in the other groups. The behaviour of $\langle J_{\phi}\rangle$ aligns with $\langle L_{z}\rangle$, as anticipated. Lastly, $\langle J_{z}\rangle$ remains generally constant across all metallicity-stratified groups, except for the two most metal-poor ones, which show higher vertical action. This may reflect either remaining interlopers from the thick disc or perturbations due to the pericentric passage of Sagittarius in the outer disc.
@@
    Finally, some results regarding the last two most metal-poor groups must be considered cautiously, due to remaining potential interlopers from the thick disc, which may not have been completely removed. However, if indeed all these stars belong to the Galactic thin disc, we identify at least three potential indicators that may suggest the influence of the pericentric passage of Sagittarius (which happened around 6 Gyr ago) on the outer disc. Group 4 and, especially, Group 3 consist of stars that:
@@
      reach significant heights above the Galactic plane, which suggests a potential influence from the pericentric passage of Sagittarius; however, it is important to acknowledge that outward churning could also contribute to this feature, meaning this vertical displacement may result from a combination of both mechanisms;
@@
      exhibit angular momentum and action components (notably $\langle J_{z}\rangle$) that differ markedly from other metallicity-stratified groups;
@@
      and display a higher proportion of stars not associated with the disc, as evidenced by the preliminary chemo-dynamic analysis. Although we do not explore these outliers in detail in the main body of the paper, their increased frequency in these groups may be an indication of a possible disturbance in the outer disc, although it is hard to differentiate them from the typical stars of the halo (usually metal-poor) at this stage.
@@
Our findings provide important insights into the migration patterns of stars across the Galaxy, revealing how metallicity, age, angular momentum, and action trace the history of their motions. By using a streamlined GAM approach to extend chemical enrichment models, we deepen our understanding of the MW’s dynamic structure and the complex forces influencing its stellar populations.

### v2

@@ 3.2.3 Other signs of migration
In the right panel, we greatly enlarged the marker sizes to make the high $\langle e\rangle$ region more prominent. As a result, stars with the largest eccentricities — least likely to be affected by churning — cluster on the left side of the plot (e.g. Khoperskov et al., 2020). These stars show a wider range of metallicities, smaller $\langle R_{\rm{g}}\rangle$ values (closer to the Galactic centre), and tend to be older. Conversely, in the left panel, stars with the largest radial differences, $\Delta R$, span a range of $\langle R_{\rm{g}}\rangle$ values but are predominantly metal-poor, with ages ranging from young to intermediate (about 0 to 4–5 Gyr). Although the marker size indicates stars with larger variation in their past and current orbital radii, which can be a sign of radial migration, it does not necessarily confirm it, since we are not accounting for errors-in-measurements here. This is discussed further in Sect. 3.3.
@@
3.3 Churned versus blurred/undisturbed stars
@@ 3.3.2 Angular momentum features of motion-stratified stellar populations
Figure 18: Median angular momenta in the $r$, $\phi$, and $z$ directions – $\langle L_{r}\rangle$, $\langle L_{\phi}\rangle$, $\langle L_{z}\rangle$ – plotted against the median metallicity, $\langle\rm{[Fe/H]}\rangle$. Each panel presents the angular momentum in a specific direction. Marker colours represent distinct HC groups, consistent with the other figures in this paper, while marker shapes indicate different stellar movements: inward churning, outward churning, and blurring (labelled as ‘equal’). The lines serve as visual guides rather than regressions, highlighting potential trends: the solid line connects HC groups with inward churning, the dashed line corresponds to outward churning, and the dotted line represents stars with unchanged radii (undisturbed or blurred).
In Fig. 18, we display $\langle L_{r}\rangle$, $\langle L_{\phi}\rangle$, and $\langle L_{z}\rangle$ against $\langle\rm{[Fe/H]}\rangle$, [Fe/H] being a rough proxy for $\langle R_{\rm{b}}\rangle$ as illustrated by the radial metallicity gradients shown in this paper. The colours in this figure align with those used throughout this paper, representing the different HC metallicity-stratified groups. Marker shapes indicate the type of movement (inward/outward-churned; and blurred/undisturbed, tagged as ‘equal’), with distinct lines connecting markers of the same movement to guide the eye towards potential trends (without performing a formal regression analysis).
@@ 3.3.4 $\langle R_{\rm{b}}\rangle$, $\langle R_{\rm{g}}\rangle$, and the OLR
Building on the discussion developed right before the beginning of Sect. 3.3.1, the bar structure likely influences stars formed in the inner regions of the MW, whereas stars moving inward may be more sensitive to the transient spiral arms than to the bar. Although there is ongoing debate regarding the length of the MW’s bar and its components (such as the so-called long-bar; López-Corredoira et al., 2007; Wegg & Gerhard, 2013; Portail et al., 2017; Lucey et al., 2023), recent estimates place its length at $\approx$ 3.5 kpc (Lucey et al., 2023), which could constrain the range of the bar’s influence, particularly in its proximity. These estimates agree with predictions from Khoperskov et al. (2020), which indicate that the bar—and more specifically, its deceleration—contributes to the churning of stars from the inner Galaxy outward towards the OLR.
@@
Figure 19: Same as Fig. 18 but for the action components ($\langle J_{r}\rangle$, $\langle J_{\phi}\rangle$, $\langle J_{z}\rangle$), instead of angular momentum.
Figure 20: $\langle R_{\rm{b}}\rangle$ against $\langle R_{\rm{g}}\rangle$ for all the metallicity- and motion-stratified groups of our sample. The general features of this plot are consistent with those observed in Figs. 18 and 19. The expected outer OLR radius (estimated by Dehnen, 2000) is marked in yellow.
@@ 3.4 Determining the probable birth radius of the Sun
Here, we investigate what our GAM estimates as the most probable location of the Sun’s $R_{\rm b}$. As input parameters, we adopted the Sun’s age, $t_{\odot}=4.775\pm 0.039$ Gyr, as provided by Bonanno & Fröhlich (2015). Specifically, we selected the final estimate from Table 3 of Bonanno & Fröhlich (2015), which the authors regard as the most reliable. This estimate was chosen based on the highest Bayes factor, derived using the Irwin+AdelR+Asplund model. For the Sun’s metallicity, we followed the convention [Fe/H]=0 but allowed for an error margin of 0.03 dex, as suggested by the non-local thermal equilibrium estimate presented by Korn et al. (2003) and further discussed in Asplund et al. (2009).
@@ 4 Summary and conclusions
In this study, we have harnessed the power of a GAM to refine the spatial and chemical grids of established chemical enrichment models for the MW (Magrini et al., 2009). Moreover, this GAM framework was adeptly employed to estimate the birth radii of a select group of Gaia-ESO stars encompassing a diverse range of metallicities spanning from metal-poor to super-metal-rich previously classified into six metallicity-stratified groups through HC. Detailed elucidation of our data curation and processing methodologies can be found in the comprehensive exposition by Dantas et al. (2023). Our overarching aim revolved around a quantitative and qualitative exploration of the dynamic effects manifesting within these stars, notably the phenomena of blurring and churning (the latter is also known as radial migration). Our conclusions are as follows:
@@
    The GAM has been proven to be a useful method for directly deriving the Galactocentric birth radii, $\langle R_{\rm{b}}\rangle$, of the stars in our sample by employing a minimalist approach that uses only [Fe/H] and the age of the Universe at the time of the stars’ birth ($t$). This simplicity was intentional, as introducing other available abundances – such as $\alpha$-elements, which are highly correlated with one another and with [Fe/H] – would not necessarily improve GAM’s performance. Instead, GAM’s effectiveness is more likely to improve if the underlying chemical enrichment models incorporate a broader range of independent information, such as neutron-capture element abundances, or better reflect the complexities of the MW’s evolutionary history, including multiple gas infall events. Additionally, models that provide a richer dataset with larger age and metallicity bins as well as smoother radial gradients would allow for more precise predictions. Conversely, including degenerate information does not enhance the performance of the GAM, further justifying our choice of a minimalist approach in this study.
@@
    We estimated the expected $\langle R_{\rm{b}}\rangle$ for the stars of our sample using chemical evolution models by Magrini et al. (2009), which were used to run a GAM to expand the reach of the original theoretical models. By comparing their $\langle R_{\rm{b}}\rangle$ with current guiding radii ($\langle R_{\rm{g}}\rangle$), we classified the stars of our sample according to their motion: inward churning, outward churning, and equal motion (undisturbed or blurred). As a consequence, we unveiled a significant migration pattern within the Galaxy:
@@
      We identified that about three-fourths of our sample have sustained churning, whereas the remaining roughly one-fourth is either undisturbed or subject to blurring.
@@
      The most metal-poor stars, in contrast, appear to have originated at larger Galactocentric distances and migrated inward towards the Galaxy’s central regions (approximately 43%), while the largest portion seem to have retained their $\langle R_{\rm{b}}\rangle$ (around 48%). It is important to note, however, that this group may still contain a small number of thick disc interlopers with metallicities that overlap with the thin disc. Despite this possible contamination, the general patterns in their motion remain consistent within this subset.
@@
      Stars with intermediate metallicities exhibit a transitional behaviour, with some migrating outward and others inward. In other words, as median metallicities decrease throughout our metallicity-stratified groups, the proportion of stars migrating outward decreases, and the number of those migrating inward increases. Additionally, the number of stars with orbits consistent with their $\langle R_{\rm{b}}\rangle$ also rises. This smooth transition is consistent throughout the groups.
@@
    There is a significant age gap between stars influenced by inward and outward churning. For instance, in Group 1 (the second most metal-rich group), this gap is approximately 5 Gyr, and it tends to persist or widen in other metallicity-stratified groups. Affected stars that have kept their orbital radii, on the other hand, exhibit intermediate ages. This age pattern remains consistent across all HC groups.
@@
    Due to the overall old ages of our blurred/undisturbed group, we doubt that most of these stars are undisturbed since we believe that time is an important factor that dictates the interactions between the stars and the Galactic structures. In other words, we reckon that most stars seem to have undergone blurring.
@@
    Although eccentricity is expected to play a relevant role for churned stars, we did not see any strong relation between eccentricity and type of movement (we do see that stars orbiting at lower $\langle R_{\rm{g}}\rangle$ have slightly larger eccentricities). The eccentricity for all the metallicity- and motion-stratified groups of stars in the sample is somewhat low, with median values ranging from 0.13 to 0.15. There is no observable trend across our HC groups.
@@
    The angular momentum ($L$) analysis results provide valuable insights into the processes that shape stellar motion within our sample. Notably, $\langle L_{z}\rangle$ revealed distinct trends, contrasting with the lack of significant patterns in $\langle L_{r}\rangle$ and $\langle L_{\phi}\rangle$. Across the sample, $\langle L_{z}\rangle$ consistently increases as metallicity decreases, irrespective of stellar motion type. This trend may reflect the chemical enrichment history of the MW, as $L_{z}$ traces patterns within the Galactic disc. Further, age differences between motion-stratified groups strengthen this interpretation. Outward-churned stars, for instance, tend to have higher $L_{z}$ and are systematically older within each metallicity-defined group than their blurred/undisturbed and inward-churned counterparts. Inward-churned stars, conversely, are consistently the youngest, exhibiting lower $L_{z}$ than their blurred/undisturbed peers, who hold intermediate positions in age and $L_{z}$. Another perspective (or interpretation) on the behaviour of $L_{z}$ considers the role of Galactic structures: outward churning and elevated $\langle L_{z}\rangle$ might result from interactions with these structures, whereas inward-churned stars, being younger, appear to ‘lose’ $\langle L_{z}\rangle$ as they move inward. These findings underscore the possible influence of Galactic structures on stellar motion types, particularly concerning $L_{z}$. However, the degree to which these ages align with the timescales required for inward or outward churning remains an open question.
@@
    The results observed in our analysis of the action components – $\langle J_{r}\rangle$, $\langle J_{\phi}\rangle$, $\langle J_{z}\rangle$ – corroborate findings from other parameters and reinforce the interpretations discussed. Specifically, $\langle J_{r}\rangle$, which quantifies radial orbital excursions as an integrated quantity, provides key insights into radial orbital stability. Our analysis suggests that stars in Group 6 may have stabilised their orbits earlier. A reasonable hypothesis is their proximity at birth to the OLR radius. This resonance, known to influence orbital dynamics, could have accelerated the damping of radial oscillations, allowing these stars to achieve orbital stability earlier compared to stars in the other groups. The behaviour of $\langle J_{\phi}\rangle$ aligns with $\langle L_{z}\rangle$, as anticipated. Lastly, $\langle J_{z}\rangle$ remains generally constant across all metallicity-stratified groups, except for the two most metal-poor ones, which show higher vertical action. This may reflect either the remaining interlopers from the thick disc or perturbations due to the pericentric passage of Sagittarius in the outer disc.
@@
    Finally, some results regarding the last two most metal-poor groups must be considered cautiously due to the interlopers that may potentially remain in the thick disc, as they may not have been completely removed. However, if all of these stars indeed belong to the Galactic thin disc, we identify at least three potential indicators that may suggest the influence of the pericentric passage of Sagittarius (which happened around 6 Gyr ago) on the outer disc. Group 4 and, especially, Group 3 consist of stars that do the following:
@@
      The stars reach significant heights above the Galactic plane, which suggests a potential influence from the pericentric passage of Sagittarius; however, it is important to acknowledge that outward churning could also contribute to this feature, meaning this vertical displacement may be from a combination of both mechanisms.
@@
      The stars exhibit angular momentum and action components (notably $\langle J_{z}\rangle$) that differ markedly from other metallicity-stratified groups.
@@
      Groups 3 and 4 display a higher proportion of stars not associated with the disc, as evidenced by the preliminary chemo-dynamic analysis. Although we did not explore these outliers in detail in the main body of the paper, their increased frequency in Groups 3 and 4 may be an indication of a possible disturbance in the outer disc. Nevertheless, it is hard to differentiate them from the typical stars of the halo (usually metal-poor) at this stage.
@@
Our findings provide important insights into the migration patterns of stars across the Galaxy, revealing how metallicity, age, angular momentum, and action trace the history of their motions. By using a streamlined GAM approach to extend chemical enrichment models, we provide the means to deepen understanding of the MW’s dynamic structure and the complex forces influencing its stellar populations.